# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 3.1 Optimal recovery of $f$
Let ${\cal F}:=U({\cal B})$ where ${\cal B}=B_{q}^{s}(L_{p}(\Omega))$ with $0\leq q,p\leq\infty$ and $s>\frac{d}{p}$. We define the tensor product grid $G_{k,r}$ of $\overline{\Omega}=[0,1]^{d}$ for $r\geq 2$,
with spacing $h=2^{-k}(r-1)^{-1}$ (the role of $r$ will be made clear below) and prove that, see Theorem 3.1,
$R^{*}({\cal F},G_{k,r})_{X}\lesssim\widetilde{m}^{-\alpha_{X}}.$
Moreover, we show that for any choice of data cites ${\cal X}\subset\overline{\Omega}$ with $|{\cal X}|=\widetilde{m}$, we have
@@
and therefore $R_{\widetilde{m}}^{*}({\cal F})_{X}\asymp\widetilde{m}^{-\alpha_{X}}$. The proof of this theorem is presented in the Appendix.
@@
    $R_{\widetilde{m}}^{*}({\cal F})_{X}\lesssim\log(\widetilde{m})\widetilde{m}^{-\alpha_{-1}},\qquad R_{\widetilde{m}}^{*}({\cal F})_{X}\gtrsim\widetilde{m}^{-\alpha_{-1}},\ \qquad\widetilde{m}\geq 1,$ (3.13)
@@
Moreover, when $\widetilde{m}=(r2^{k})^{d}$, the upper bounds in each of these estimates are obtained when we take the data sites ${\cal X}=G_{k,r}$, provided $r>\max(s,1)$. In this case, the approximant to $f$ is given by the continuous piecewise polynomial
@@
described in (2.16), where ${\cal T}_{k}={\cal T}_{k}(\Omega)=\bigcup T$ is the collection of all simplicies arising from the (Kuhn-Tucker) conforming decomposition of the dyadic cubes in ${\cal D}_{k}(\Omega)$, the $\chi_{T}$’s are the characteristic functions of $T$, and the $L_{T}(f)$’s are the polynomials of order $r$ ($L_{T}(f)\in{\cal P}^{d}_{r}$), gotten by interpolating $f$ at the data points $T\cap G_{k,r}$.
@@
$\|f\|_{H^{-1}(\Omega)}\leq\begin{cases}C(d)\|f\|_{L_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\delta}(\Omega)},\quad d\geq 3,\\ \\ \frac{C}{\tau-1}\|f\|_{L_{\tau}(\Omega)},\quad\,\,d=2,\quad\tau>1,\end{cases}$ (3.14)
@@ 3.2 Optimal recovery of $g$
In this section, we study the optimal recovery rate of functions $g\in{\cal G}$ in the norm of $H^{1/2}(\partial\Omega)$. Let $Tr=Tr_{\partial\Omega}$ be the trace operator onto $\partial\Omega$ and let ¯$\overline{\cal B}:=B_{\bar{q}}^{\bar{s}}(L_{\bar{p}}(\Omega))$, where we fix $\bar{s},\bar{p},\bar{q}$ and assume that $\overline{s}>d/\overline{p}$ so that we are sure that the functions in $\overline{\cal B}$ are continuous on $\overline{\Omega}$. We consider the case $0<\bar{p}\leq 2$, which together with the restriction $d\geq 2$ implies that $\bar{s}-d(1/\bar{p}-1/2)>1$, and thus the unit ball of $\overline{\cal B}$ compactly embeds into $H^{1}(\Omega)$. We define the model class
@@
denote the set of those grid points of $G_{k,r}$ that are on the boundary $\partial\Omega$. We will use the points ${\cal Z}=\overline{G}_{k,r}$ to recover $g$. Note that the number $\bar{m}$ of data sites in $\overline{G}_{k,r}$ is
@@
$Tr(S_{k}^{*}(v))=\overline{S}_{k}(g).$
This follows from the fact that each Lagrange interpolant $\phi_{i}$ centered at ${\bf x}_{i}\in G_{k,r}\setminus\overline{G}_{k,r}$ vanishes at the faces of the simplex from the corresponding Kuhn-Tucker decomposition that contains ${\bf x}_{i}$. We also want to point out that $\overline{S}_{k}(g)$ does not depend on $v$.
The proof of the following result is provided in the appendix.
@@
Let ¯$\overline{\cal B}=B_{\overline{q}}^{\overline{s}}(L_{\bar{p}}(\Omega))$ with $\bar{s}>d/\bar{p}$, $0<\bar{p}\leq 2$, $0<q\leq\infty$, and let ${\cal G}=\{Tr(v):\|v\|_{\overline{\cal B}}\leq 1\}$. Then the optimal recovery rates of the model classe ${\cal G}$ in $H^{1/2}(\partial\Omega)$ is
@@
with constants of equivalency independent of $\overline{m}$.
Moreover, when $\overline{m}=\#(\overline{G}_{k,r})$, the upper bound in (3.19) is achieved when we take the data sites ${\cal Z}=\overline{G}_{k,r}$, provided $r>\max(s,1)$ and the approximant to $g$ is given by the function $\overline{S}_{k}(g)$.
To understand the exponent $\beta$ in (3.19) better, let $v$ be a function in $U(B_{q}^{\bar{s}}(L_{\bar{p}}(\Omega)))$ whose trace is $g$. Note that $\beta$ can be rewritten as
$\beta=\frac{\overline{s}-1-(\frac{d}{\bar{p}}-\frac{d}{2})}{d-1},$
where the numerator is the excess regularity of $v$ in $H^{1}(\Omega)$ and thereby corresponds to the excess regularity of $g$ in $H^{1/2}(\partial\Omega)$, and the denominator is $d-1$ since $\partial\Omega$ has dimension $d-1$.
@@ 3.3 Optimal recovery of $u$
under a given budget $m=\widetilde{m}+\overline{m}$ of data observations. Recall that ${\cal F}:=U(B_{q}^{s}(L_{p}(\Omega)))$, where $s>d/p$ and $0<q,p\leq\infty$, ${\cal G}:=Tr(U(B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega)))$, where $\overline{s}>d/\overline{p}$ and $0<\overline{p}\leq 2$, and
@@
where $\alpha:=\alpha_{-1}$ is given by (3.12), $\beta$ is given by (3.19), and all constants of equivalence are independent of $m$.
@@
$\|u_{1}-u_{2}\|_{H^{1}(\Omega)}\asymp\|f_{1}-f_{2}\|_{H^{-1}(\Omega)}+\|g_{1}-g_{2}\|_{H^{1/2}(\Omega)}.$
It follows that
@@ 3.4 Final observations on optimal recovery
This section gave the optimal recovery rates for the model classes ${\cal F},{\cal G},{\cal U}$ from $\widetilde{m},\overline{m},m$, data observations, respectively. While we began by considering general model class assumptions ${\cal F}$ and ${\cal G}$ for $f$ and $g$ to be the unit balls of Besov spaces $B_{q}^{s}(L_{p})$, we found that the optimal recovery rate is the same for many of these model classes. This means that some of the model classes are superflous in that they are contained in a larger model class with the same recovery rate. It makes sense only to use these largest model classes in going forward.
Consider, for example, the model classes ${\cal F}=U({\cal B})$ with ${\cal B}=B_{q}^{s}(L_{p}(\Omega))$ for $f$. We found that in the case $d\geq 3$ and $0<p\leq\infty$, or $d=2$ and $1<p\leq\infty$, the optimal recovery rate of ${\cal F}$ in $H^{-1}(\Omega)$ is the same as its optimal recovery rate in $L_{\delta}(\Omega)$ with ${\delta}=\frac{2d}{d+2}$. Moreover, the model class ${\cal F}$ has an optimal recovery rate $m^{-\alpha/d}$ in $H^{-1}(\Omega)$ if and only if it is contained in $U(B_{\infty}^{\alpha}(L_{\gamma}(\Omega)))$ and, in addition, ${\cal F}$ embeds into $C(\Omega)$. Largest Model Classes for $f$:
@@
and any Besov model class that gives the optimal recovery rate $O(m^{-s/d})$ is contained in one of these model classes. In each of these cases, the number $s$ represents the smoothness of $f\in{\cal F}$ in $L_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\delta}(\Omega)$. These largest model classes are pictured in Figure 3.1.
@@
Largest Model Classes for $g$: We have considered the model classes ${\cal G}=Tr(B^{\bar{s}}_{\infty}(L_{\bar{p}}(\Omega))$ with $0<\bar{p}\leq 2$ and $\bar{s}>\max\{\frac{d}{\bar{p}},1\}$. Thus, all of the model classes which give a given rate $O(\overline{m}^{-\alpha})$ will all be contained in one model class
@@ 4 Numerical algorithms based on linear approximation
The above analysis of optimal recovery tells us to what extent the information $({\bf f},{\bf g})$ determines $u$. That is, the optimal recovery analysis tells us that if two functions $u,\widetilde{u}$ from our model class ${\cal U}$ satisfy the same data, then they are close in the $H^{1}(\Omega)$ norm, that is
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim R_{m}^{*}({\cal U})_{H^{1}(\Omega)}.$
However, it does not give a numerical algorithm that takes the data and creates an approximation $\widehat{u}$ to $u$ with accuracy of the optimal recovery rate. In this section, we discuss two numerical methods which would accomplish the latter task. The methods discussed in this section are not in the form of PINNs. The latter will be discussed in §9.
We assume in this section that the data sites form a tensor product grid $G$ of $\overline{\Omega}$. We know such data sites $G$ provide a near optimal recovery. We have shown that if we use the data ${\bf f}$ at $G$, then we can numerically construct a piecewise polynomial interpolant $\widehat{f}$ to $f$ at the data sites which is a near optimal (in terms of the budget $\widetilde{m}$) approximation to $f$ in the $H^{-1}(\Omega)$ norm. Similarly, we can numerically construct a $\widehat{g}$ from the data ${\bf g}$ which is a near optimal recovery of $g$ in the $H^{1/2}(\partial\Omega)$ norm.
We define $\widetilde{u}$ as the solution to (1) with right hand side $\widehat{f}$ and boundary value $\widehat{g}$. We know that $\widetilde{u}$ is a near optimal recovery of $u$ in the $H^{1}(\Omega)$ norm. Then, in view of (1.3), we have
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim R_{m}^{*}({\cal U})_{H^{1}(\Omega)}.$ (4.1)

### v2

@@ 3.1 Optimal recovery of $f$
Let ${\cal F}:=U({\cal B})$ where ${\cal B}=B_{q}^{s}(L_{p}(\Omega))$ with $0\leq q,p\leq\infty$ and $s>\frac{d}{p}$. Recall, see (2.15), the tensor product grid $G_{k,r}$ of $\overline{\Omega}=[0,1]^{d}$ for $r\geq 2$, with spacing $h=2^{-k}(r-1)^{-1}$ (the role of $r$ will be made clear below). We will prove that, see Theorem 3.1,
$R^{*}({\cal F},G_{k,r})_{X}\lesssim\widetilde{m}^{-\alpha_{X}},\quad\widetilde{m}=(r2^{k})^{d}.$
Moreover, we show that for any choice of data sites ${\cal X}\subset\overline{\Omega}$ with $|{\cal X}|=\widetilde{m}$, we have
@@
and therefore $R_{\widetilde{m}}^{*}({\cal F})_{X}\asymp\widetilde{m}^{-\alpha_{X}}$. The proof of this theorem is presented in the Appendix, §12.6.1.
@@
    $\widetilde{m}^{-\alpha_{-1}}\lesssim R_{\widetilde{m}}^{*}({\cal F})_{X}\lesssim\log(\widetilde{m})\widetilde{m}^{-\alpha_{-1}},\qquad\qquad\widetilde{m}\geq 1,$ (3.12)
@@
Moreover, when $\widetilde{m}=(r2^{k})^{d}$, the upper bounds in each of these estimates are obtained when we take the data sites ${\cal X}=G_{k,r}$, provided $r>\max(s,1)$. In this case, the approximant to $f$ is given by the continuous extension of the piecewise polynomial
@@
described in (2.18), where ${\cal T}_{k}={\cal T}_{k}(\Omega)=\bigcup T$ is the collection of all simplicies arising from the (Kuhn-Tucker) decomposition of the dyadic cubes in ${\cal D}_{k}(\Omega)$, the $\chi_{T}$’s are the characteristic functions of $T$, and the $L_{T}(f)$’s are the polynomials of order $r$ ($L_{T}(f)\in{\cal P}^{d}_{r}$) gotten by interpolating $f$ at the data points $\overline{T}\cap G_{k,r}$.
@@
$\|f\|_{H^{-1}(\Omega)}\leq\begin{cases}C(d)\|f\|_{L_{\delta}(\Omega)},\quad d\geq 3,\quad\hbox{where}\,\,\delta=\frac{2d}{d+2},\\ \\ \frac{C}{\tau-1}\|f\|_{L_{\tau}(\Omega)},\quad\,\,d=2,\quad\tau>1,\end{cases}$ (3.13)
@@ 3.2 Optimal recovery of $g$
In this section, we study the optimal recovery rate of functions $g\in{\cal G}$ in the norm of $H^{1/2}(\partial\Omega)$. Let $Tr=Tr_{\partial\Omega}$ be the trace operator onto $\partial\Omega$ and let ¯$\overline{\cal B}:=B_{\bar{q}}^{\bar{s}}(L_{\bar{p}}(\Omega))$, where we fix $\bar{s},\bar{p},\bar{q}$ and assume that $\overline{s}>d/\overline{p}$ so that we are sure that the functions in $\overline{\cal B}$ are continuous on $\overline{\Omega}$. We consider the case $0<\bar{p}\leq 2$, which together with the restriction $d\geq 2$ implies that $\bar{s}-d(1/\bar{p}-1/2)>1$, and thus the unit ball of $\overline{\cal B}$ compactly embeds into $H^{1}(\Omega)$.
We define the model class
@@
denote the set of those grid points of $G_{k,r}$ that are on the boundary $\partial\Omega$. We will use the points ${\cal Z}=\overline{G}_{k,r}$ to recover $g$. Note that the number $\overline{m}$ of data sites in $\overline{G}_{k,r}$ is
@@
$Tr(S_{k}^{*}(v))=\overline{S}_{k}(g).$ (3.18)
This follows from the fact that each Lagrange interpolant $\phi_{i}$ centered at ${\bf x}_{i}\in G_{k,r}\setminus\overline{G}_{k,r}$ vanishes at the faces of the simplex from the corresponding Kuhn-Tucker decomposition that contains ${\bf x}_{i}$. Note that $\overline{S}_{k}(g)$ does not depend on $v$. According to Theorem 2.2, $S_{k}^{*}(v)\in H^{1}(\Omega)$ since $v$ belongs to unit ball of $\overline{\cal B}$. Then, (3.18) gives that $S_{k}(g)\in H^{1/2}(\partial\Omega)$.
The proof of the following result is provided in the Appendix, §12.6.2.
@@
Let ¯$\overline{\cal B}=B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega))$ with $\overline{s}>d/\bar{p}$, $0<\overline{p}\leq 2$, $0<\overline{q}\leq\infty$, and let ${\cal G}=\{Tr(v):\|v\|_{\overline{\cal B}}\leq 1\}$. Then the optimal recovery rates of the model class ${\cal G}$ in $H^{1/2}(\partial\Omega)$ is
@@
with constants of equivalency independent of $\overline{m}$. Moreover, when $\overline{m}=\#(\overline{G}_{k,r})$, the upper bound in (3.19) is achieved when we take the data sites ${\cal Z}=\overline{G}_{k,r}$, provided $r>\max(s,1)$ and the approximant to $g$ is given by the function $\overline{S}_{k}(g)$.
To understand the exponent $\beta$ in (3.19) better, let $v$ be a function in $U(B_{q}^{\bar{s}}(L_{\bar{p}}(\Omega)))$ whose trace is $g$, and rewrite $\beta$ as
$\beta=\frac{\overline{s}-1-(\frac{d}{\bar{p}}-\frac{d}{2})}{d-1}.$
Here the numerator is the excess regularity of $v$ in $H^{1}(\Omega)$ and thereby corresponds to the excess regularity of $g$ in $H^{1/2}(\partial\Omega)$, and the denominator is $d-1$ since $\partial\Omega$ has dimension $d-1$.
@@ 3.3 Optimal recovery of $u$
under a given total budget $m=\widetilde{m}+\overline{m}$ of data observations. Recall that ${\cal F}:=U(B_{q}^{s}(L_{p}(\Omega)))$, where $s>d/p$ and $0<q,p\leq\infty$, ${\cal G}:=Tr(U(B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega)))$, where $\overline{s}>\frac{d}{\bar{p}}\geq 1$ since $d\geq 2$ and $0<\overline{p}\leq 2$, $0<\overline{q}\leq\infty$, and
@@
where
$\alpha=\frac{s}{d}-\left[\frac{1}{p}-\left(\frac{1}{2}+\frac{1}{d}\right)\right]_{+},\quad\beta=\frac{\overline{s}-1-(\frac{d}{\bar{p}}-\frac{d}{2})}{d-1},$ (3.23)
and all constants of equivalence are independent of $m$.
@@
$\|u_{1}-u_{2}\|_{H^{1}(\Omega)}\asymp\|f_{1}-f_{2}\|_{H^{-1}(\Omega)}+\|g_{1}-g_{2}\|_{H^{1/2}(\Omega)},$
and therefore
@@
Remark 3.4.
Consider $\Omega=(0,1)^{d}$ and ${\cal F}$, ${\cal G}$, ${\cal U}$ as above. Let us assume that we have a fixed budget of $\widetilde{m}$ data sites ${\cal X}$ from $\overline{\Omega}$ and a fixed number $\overline{m}$ of data sites ${\cal Z}$ on the boundary $\partial\Omega$ and denote by
$R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}:=\inf_{{\cal X}\subset\overline{\Omega},\,{\cal Z}\subset\partial\Omega,\,|{\cal X}|=\widetilde{m},\,|{\cal Z}|=\overline{m}}R^{*}({\cal U},{\cal X},{\cal Z})_{H^{1}(\Omega)}.$
It follows from (3.24) that
$R_{\widetilde{m}}^{*}({\cal F})_{H^{-1}(\Omega)}+R^{*}_{\overline{m}}({\cal G})_{H^{1/2}(\partial\Omega)}\asymp R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}.$ (3.25)
Theorem 3.1 and Theorem 3.2 then yield
$\widetilde{m}^{-\alpha}+\overline{m}^{-\beta}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}\lesssim\begin{cases}\log(\widetilde{m})\widetilde{m}^{-\alpha}+\overline{m}^{-\beta},\quad d=2,\,0<p\leq 1,\\ \widetilde{m}^{-\alpha}+\overline{m}^{-\beta},\quad\quad\quad\quad d=2,\,1<p\leq\infty\quad\hbox{or}\quad d\geq 3,\end{cases}$ (3.26)
where $\alpha,\beta$ are as in (3.23).
Now suppose instead that we are given a fixed total budget $m$ and $d\geq 2$. Then we can obtain the optimal rate $R^{*}_{m}({\cal U})_{H^{1}(\Omega)}$ by choosing $\widetilde{m}\asymp\overline{m}\asymp m/2$.
@@ 3.4 Final observations on optimal recovery
This section gave the optimal recovery rates for the model classes ${\cal F},{\cal G},{\cal U}$ from $\widetilde{m},\overline{m},m$, data observations, respectively. While we began by considering general model class assumptions ${\cal F}$ and ${\cal G}$ for $f$ and $g$ to be the unit balls of Besov spaces, we found that the optimal recovery rate is the same for many of these model classes. This means that some of the model classes are superflous in that they are contained in a larger model class with the same recovery rate. It makes sense only to use these largest model classes in going forward.
Consider, for example, the model classes ${\cal F}=U({\cal B})$ with ${\cal B}=B_{q}^{s}(L_{p}(\Omega))$ for $f$. We found that in the case $d\geq 3$ and $0<p\leq\infty$, or $d=2$ and $1<p\leq\infty$, the optimal recovery rate of ${\cal F}$ in $H^{-1}(\Omega)$ is the same as its optimal recovery rate in $L_{\delta}(\Omega)$ with ${\delta}=\frac{2d}{d+2}$. Moreover, the model class ${\cal F}$ has an optimal recovery rate $m^{-\alpha/d}$ in $H^{-1}(\Omega)$ if and only if it is contained in $U(B_{\infty}^{\alpha}(L_{\delta}(\Omega)))$ and, in addition, ${\cal F}$ embeds into $C(\overline{\Omega})$.
Largest Model Classes for $f$:
@@
and any Besov model class that gives the optimal recovery rate $O(m^{-s/d})$ is contained in one of these model classes. In each of these cases, the number $s$ represents the smoothness of $f\in{\cal F}$ in $L_{\delta}(\Omega)$. These largest model classes are pictured in Figure 3.1.
@@
Largest Model Classes for $g$: We have considered the model classes ${\cal G}=Tr(B^{\bar{s}}_{\infty}(L_{\bar{p}}(\Omega))$ with $0<\bar{p}\leq 2$, $d\geq 2$, and $\bar{s}>\max\left(\frac{d}{\bar{p}},1\right)$. Thus, the model classes which give a rate $O(\overline{m}^{-\alpha})$ will all be contained in one model class
@@ 4 Numerical algorithms based on linear approximation
Recall that we consider the case when $\Omega=(0,1)^{d}$, $d\geq 2$, ${\cal F}:=U({\cal B})$, where ${\cal B}=B_{q}^{s}(L_{p}(\Omega))$ with $0\leq q,p\leq\infty$ and $s>\frac{d}{p}$, and ${\cal G}=\{Tr(v):\|v\|_{\overline{\cal B}}\leq 1\}$, where ¯$\overline{\cal B}=B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega))$ with $\overline{s}>d/\bar{p}$, $0<\overline{p}\leq 2$, $0<\overline{q}\leq\infty$. For these right hand sides and boundary values, we consider the solution set
[…]
The above analysis of optimal recovery tells us to what extent the information $({\bf f},{\bf g})$ determines $u$. That is, the optimal recovery analysis tells us that if two functions $u,u^{\prime}\in{\cal U}$ and satisfy the same data, namely $u,u^{\prime}\in{\cal U}_{\rm data}$, then they are close in the $H^{1}(\Omega)$ norm. More precisely, if $u^{*}$ is the center of a Chebyshev ball $B({\cal U}_{\rm data})$, then
$\|u-u^{\prime}\|_{H^{1}(\Omega)}\leq\|u-u^{*}\|_{H^{1}(\Omega)}+\|u^{*}-u^{\prime}\|_{H^{1}(\Omega)}\leq 2R_{m}^{*}({\cal U})_{H^{1}(\Omega)}.$ (4.1)
However, our analysis does not give a numerical algorithm that takes the data and creates an approximation $\widehat{u}$ to $u$ with accuracy of the optimal recovery rate. In this section, we discuss two numerical methods which would accomplish the latter task. The methods discussed in this section are not in the form of PINNs. The latter will be discussed in §8.
We assume in this section that the data sites form tensor product grids, that is, ${\cal X}=G_{k,r}$ and ${\cal Z}=\overline{G}_{\bar{k},r}:=G_{\bar{k},r}\cap\partial\Omega$. We know such data sites provide a near optimal recovery. We have shown that if we use the data ${\bf f}$ at $G_{k,r}$, then we can numerically construct a continuous piecewise polynomial interpolant $\widehat{f}$ to $f$ at the data sites which is a near optimal (in terms of the budget $\widetilde{m}=(r2^{k})^{d}$) approximation to $f$ in the $H^{-1}(\Omega)$ norm. Similarly, we can numerically construct a $\widehat{g}$ from the data ${\bf g}$ which is a near optimal recovery of $g$ in the $H^{1/2}(\partial\Omega)$ norm in terms of the budget $\overline{m}\asymp 2^{\bar{k}(d-1)}$. The following lemma holds.
Lemma 4.1.
Let $\widetilde{u}$ be the solution to (1) with right hand side $\widehat{f}:=S^{*}_{k}(f)$ and boundary value $\widehat{g}:=\overline{S}_{k}(g)$. Then
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)},\quad\hbox{where}\quad\widetilde{m}\asymp 2^{kd},\quad\overline{m}\asymp 2^{\bar{k}(d-1)}.$ (4.2)
If instead, we are given a fixed total budget $m$, then
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim R_{m}^{*}({\cal U})_{H^{1}(\Omega)},\quad\hbox{provided}\quad\widetilde{m}\asymp\overline{m}\asymp m/2.$ (4.3)
Proof.
We have that $\widehat{f}:=S^{*}_{k}(f)\in C(\overline{\Omega})\cap H^{1}(\Omega)$, see Theorem 3.1, and $\widehat{g}:=\overline{S}_{k}(g)\in H^{1/2}(\partial\Omega)\cap C(\partial\Omega)$, see Theorem 3.2. Note that we cannot use (4.1) to bound $\|u-\tilde{u}\|_{H^{1}(\Omega)}$ since $\hat{f}\notin{\cal F}$ and $\hat{g}\notin{\cal G}$ (they do not have the necessary smoothness) and therefore $\tilde{u}\notin{\cal U}$. In this case, in view of (1.4) with $v=\widetilde{u}$, we have
$\|u-\widetilde{u}\|_{H^{1}(\Omega)}\lesssim\|f-\hat{f}\|_{H^{-1}(\Omega)}+\|g-\hat{g}\|_{H^{1/2}(\partial\Omega)}\lesssim R_{\widetilde{m}}^{*}({\cal F})_{H^{-1}(\Omega)}+R^{*}_{\overline{m}}({\cal G})_{H^{1/2}(\partial\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)},$
where we have used (3.25). This proves (4.2). In view of Remark 3.4, given $m$ data sites, the solution $\widetilde{u}$ that corresponds to the choice of $\widetilde{m}\asymp m/2$ points on $\overline{\Omega}$ and $\overline{m}\asymp m/2$ points on $\partial\Omega$ satisfies (4.3) and thus is a near optimal recovery of $u$ in the $H^{1}(\Omega)$ norm. ∎